OpenStudy (anonymous):

Use the quadratic formula: 3-8v-5v^2=2v Help!!!

OpenStudy (anonymous):

Show work? Thanks :)

OpenStudy (anonymous):

5v^2+10v-3=0 v1,2 = (-10 +/- (10^2 - 4*5*(-3))^0.5) / 10 v1,2 = (-10 +/- (100 + 60)^0.5) / 10 v1,2 = (-10 +/- 12.649) / 10 v1,2 = 0.2649, -2.2649

OpenStudy (anonymous):

Thanks! :D Can you check my answer for this problem? Same directions. -3y^2=6y-10 My answer was -1+/-2 sqrt. 39/-6

OpenStudy (anonymous):

im sorry i dont understand your answer

OpenStudy (anonymous):

ok ok im sorry too sleepy

OpenStudy (anonymous):

it's alright, i think i got it :D thank for ur help! :D

OpenStudy (anonymous):

as long as the /-6 is only for the square root you are right :)